OpenStudy (anonymous):

A 10-foot ladder leans against a wall with its foot braced 3 feet from wall’s base. How far up the wall does the ladder reach?

OpenStudy (anonymous):

 so the height is=sqrt(10^2-3^2). evaluate yourself.

OpenStudy (pratyush5):

height = sqrt(100 - 9 ) = sqrt (91) approx 9.53